OpenStudy (anonymous):

For which value of x is f (x) = -65 given f (x) = -3x2 - 2x For which value of x is f (x) = -65 given f (x) = -3x2 - 2x @Mathematics

OpenStudy (anonymous):

29.5

OpenStudy (anonymous):

set -3x^2-2x=-65 3x^2+2x-65=0 set equal to 0 (x+5)(3x-13)=0 factor x=5=0 or 3x-13=0 zero-factor theorem x=-5 or x=13/3 solve